A sales person has 7 products that he wishes to display at a national convention. He can display only 4. The order in which he displays the products is immaterial. How many displays does he have to choose from?
Answer by fractalier(6550)

Seven chosen four at a time is 7C4 and equals
7! / (7-4)!(4!) = 7*6*5 / 3*2*1 = 35 possible displays